A curve is such that the length of the polar radius of any pt. on the curve is equal to length of tangent at this pt.. Form diff. eqn. and find the curve ?

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mayank

let a point on the curve is (x,y)

length of polar radius  r = √x2+y2

length of tangent  = [y√{1+(dy/dx)2} ]/(dy/dx) 

so   [y√{1+(dy/dx)2} ]/(dy/dx)  = √x2+y2

     square and simplyfy

  y=±xdy/dx

dx/x = ±dy/dx

intigrate

 lnx  =± lny  +lnC

+ve sign

x=Cy

-ve sign

x=C/y

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin